HUBBLE SPACE TELESCOPE DAILY REPORT #2315 PERIOD COVERED: 0000Z (UTC) 02/23/99 - 0000Z (UTC) 02/24/99 Daily Status Report as of 055/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.3 Completed STIS/MA1 7298 (A Snapshot Survey for UV bright Quasars at redshift below 3) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet snapshot observations of the bright quasar CS0261 which has a redshift (Z) below 3.0. The goal is to find quasars bright enough in the ultraviolet for spectroscopic follow-up in the range of 900-1200 Angstroms. A wave length calibration image was also taken. The observation were completed as scheduled, and no anomalies were reported. 1.4 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 6469 (Pyramid Imaging of Circumstellar Material About Nearby Stars) The WF/PC-2 was used to make observations of the curcumstellar material around the nearby stars HD22049 and HD22049. The WF/PC-2 pyramid was used as a coronagraph to detect the circumstellar material. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/MA2 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD/MA1 7937 (MAMA Off Axis Sensitivity {Vignetting}) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of G191B2B in order to study off axis vignetting in the MA1 detector. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.8 Completed WF/PC-2 6357 (The Nuclear Morphology of Elliptical Galaxies) The WF/PC-2 was used to make images of the early type, large, elliptical galaxy NGC-2699 (taken from a list of 150 galaxies) in order to study its nuclear morphology. This survey will complement Illingworth's snapshot study of later Hubble types. The observations were completed as scheduled, and no anomalies were reported. 1.9 Completed Five Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.10 Completed Four Sets of FOC/96 6350 (FOC spectroscopy of BALQSOs) The Faint Object Camera (f/96) was used to make ultraviolet spectrographic images of BALQSOs 1303+3048, 0932+5006, 1336+1335 and 1208+1535 which are known to show broad emission lines from abundant ions. This will help us improve our knowledge of these BAL class objects. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 8022 (Supplement to the STIS Dynamics IKP) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of galaxy NGC-4552. Detection of supermassive black holes in normal galaxies is one of the grand challenges that HST was designed to meet. A wavelength calibration observation and a CCD flat were also completed. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed STIS/CCD/MA1 7432 (Morphological and Spectral Evolution of Galaxies) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations to study the morphological and spectral evolution of the galaxy 3C63-1. A library of ultraviolet images of normal galaxies at low redshifts is essential to our understanding of the properties and evolution of the high redshift universe. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared twice. CCS string "A" was configured for real-time operations at 054/1400Z. A TTR was filed for data hits occurring at 055/002731Z that affected 11 minor frames. At 055/083701Z, as the quality-checked TDE ephemeris table was being loaded and dumped, all spacecraft telemetry was lost. Subsequently, bad (high) on-board HGA-1 Y-axis torque counts were observed, and body rate error tests failed. The HST entered inertial hold safemode at 055/093737Z. Although, ephemeris loads had been successfully built, checked, uplinked, and dumped using CCS, the decision was made to return to PRS real-time support at 055/0957Z. Observations occurring during the period of the next report were affected. The Gyro-3 drift bias trend shows a decrease since the last report. The rate of change over the last two days is 34.49 arcseconds/hour/day (was 38.94). This trend still exceeds the limit of 10.8 arcseconds/hour/day for Gyro-3 reinstallation into the control loop. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH